The sum of Cathy and Tony's age is 65. Eight years ago Cathy's age was 2/5 of Tony's age. How old are they now?
Answer by richwmiller(17219)

c+t=65,
c-8=2/5(t-8)
c = 22, t = 43
check
14=2/5*(35)
14=2*7
ok